A heat engine is supplied with 278 k3/s of heat at a constant fixed temperature of 283°C and the best rejection takes place at 5°C The following results were reported (i) 208 klis are rejected, (ii) 139 klis are rejected, and (ii) 70 kJ/s are rejected Classify which of the results report a reversible cycle or irreversible cycle or impossible results

1
Expert's answer
2022-01-25T14:07:45-0500

1.

"\Sigma\frac QT=\frac{278}{556}-\frac{208}{278}=-0.248<0," cycle is irreversible.

2.

"\Sigma\frac QT=\frac{278}{556}-\frac{139}{278}=0," cycle is reversible.

3.

"\Sigma\frac QT=\frac{278}{556}-\frac{70}{278}=0.248>0," this cycle is impossible by second law of thermodynamics, i.e. Clausius inequality.